OpenStudy (anonymous):

Evaluate 5^6 divided by 5^3

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

5^6 = 5.5.5.5.5.5 5^3 = 5.5.5 \[\frac{5.5.5.5.5.5}{5.5.5}\] \[\frac{5.5.5\cancel{5.5.5}}{\cancel{5.5.5}}\]

OpenStudy (hba):

Lets see it this way, \[\frac{ 5*5*5*5*5*5 }{ 5*5*5 }\]

OpenStudy (hba):

Oh @amistre64 has already done it for you.

OpenStudy (anonymous):

thanks anyways